OpenStudy (anonymous):

Suppose f(x) = x^2 - 3 x - 1. Find the following information for this function. The coordinates of the vertex are? This parabola is what? concave up,down or both or neither

OpenStudy (anonymous):

vertex ( 3/2, -1) concave up - remember that F(x)= ax^2 + bx + c. concave direction is based on the sign of A - if negative: concave down and if positive : concave up

OpenStudy (anonymous):

the y intercepts occur at?

OpenStudy (anonymous):

FIND THE axis of symmetry by x = -b/2a vertex is -b/2a, f(-b/2a)

OpenStudy (anonymous):

wait I now its a parabola and it contains a minimum but how do I find that?

OpenStudy (anonymous):

min is calculatd by y >= f(-b/2a)

OpenStudy (anonymous):

ok so what is b and a?

OpenStudy (anonymous):

3/2 and -1?

OpenStudy (anonymous):

F(x)= ax^2 + bx + c standard f(x) = x^2 - 3 x - 1 your problem so a = 1 b = -3 c = -1

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

quadratic equation right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

my son is learning this now and i just taught him - that is why i can answer so quick

OpenStudy (anonymous):

oh ok